MP3 meeting 1 July 2009
Minutes of last meeting
A few changes were integrated in the minutes of the last meeting
Report from the last days of Powering test on S2-3:
David reported on the past weekend Powering Tests: to be reported that the 2008 Analysis tools seem to work properly.
The 2009 analysis Tools are visible on the computers but not accessible yet. Test (on 600 and 120 A correctors) very quiet.
Arjan had nothing special to report from Powering Test on Monday.
Jean-Philippe reported about Tuesday Powering Tests:
As also reported in last Matteo’s Status Report (e-mail of Tueday 30 June 6:10 PM):
- Few 600 A correctors (in triplets) show quenches during rump-up; they were OK at 2008 Powering Tests.
- Orbit corrector circuits PNO.a3 analysis: in the Powering Procedure is asked to check for Udiff value, but in order to check this value it must be launched the PCS automatic analysis that seems not logic. (ACTION: Adriaan, Walter)
- Nuria and David reported that the post mortem file generated by the detector of the RDS.A23B1 had more than half an hour delay (also in the timestamps). The signal shows a trip and not a quench of the magnet. To be followed.
- 120 correctors in the inner triplet. One of the signals used for the analysis is coming only to the FGC port mortem but it is not logged in the database. Nuria will contact Stephen Page to see if the signal can be recorded with the others. (After the meeting, it was checked that there is no room in the bus for this signal. Another way to see the voltage of the leads during the test needs to be found)
For other points refer to Matteo’s Status Report.
Knud presentation on proposed modification of Energy Extraction system for RB and RQ:
Request was done by Lyn to evaluate how to reduce the decay time of the RB and RQ circuits by increasing the energy extraction resistance. The point was also discussed yesterday at the TE-TM Meeting.
NOTE: ALL WHAT PROPOSED APPLY ONLY FOR I≤ 6850 A (i.e. 4 TeV)!
RB Circuit: Knud presented 3 proposals in order to decrease the tau of the circuit. The maximum current at which this system has proven to work decreases proportionally to tau. All three solutions are compatible with 91 A/s maximum current decay at which all sectors have been tested. The voltage across the circuit is not an issue. Arjan added that higher de-ramp rates were seen in sector 5-6. This limit can thus be removed but needs to be checked during hardware commissioning. The 3 proposals can very easily implemented by appropriate switching on the 3 dump resistors connections. The first solution was already in place for sector 2-3 during last weekend powering tests.
RQ Circuits: configuration of the resistor is less easy to be changed, but it can be done. With the proposed modification di/dt will pass from 310 to 457 A/s. This should be not a problem for the MQ (test in SM18 last month showed no problem of quench back up to 1000 A/s). Tau will change from 30 to ~15 s, V to be assessed.
Knud informed the committee about a new request to delay the switch opening between the two switches of the dipole EE circuit. Today, they are delayed of 330 ms from the PIC signal to allow the beam extraction to the dump. It is proposed to add ~ 500 ms more in order to limit the initial V rising.
- Gianluigi comment: it must be well CLARIFY which is the next PNO for RB and RQ : is it 4 or 5 TeV equivalent? It seems that the two values are actually utilized by different Teams.
AOB:
- Nuria reported about case of MB2413 in S1-2 that showed 3 QH (HF) circuits with lower insulation level to ground and to coil. Additional diagnostics were done in this magnet to asses the integrity of the remaining LF quench heaters. The insulation fault on the HF QH was re-confirmed. The cool-down of the sector started on Friday and the next step will be to repeat the tests at cold. If the problem is confirmed, a safe configuration with LF heaters is possible and will be prepared. A Roxie simulation of this configuration can be done in the following weeks.
The magnet was not showing any problem at assembly (checked by Michele) and at Cold Test (checked by Marta/Nuria).
- Also reported about QH problem in MB1263 (S. 6-7, quench heater circuit open). The problem was finally identified at the level of the IFS box and repaired. The magnet must be reconfigured with the standard QH configuration (ACTION: KNUD).
- Bob gave some pre-presentation about the analysis of the faulty splices in S.4-5. In this Sector we are probably observing the worst cases of high resistance of splices (higher is up to today ~ 70 microOhm).
A campaign of measurements at 80 and 300 K was done on RB and RQ circuits (worst case is MBA B16-C15.L5 interconnection). The splice in question was then open and shows an extremely poor execution quality.
About 80K/300K measurement correlation: it is partially present for RB splices, not at all present for RQ splices. A full presentation on the subject is planned at the next MP3 Meeting.
M. Modena, N. Catalan-Lasheras 1 July 2009
There are minutes attached to this event.
Show them.